OpenStudy (anonymous):

Use the formula for simple interest, I = P × r × t, to find the missing quantity. I = $960; P = $6,000; r = 8%

OpenStudy (texaschic101):

I = PRT I = 960 P = 6000 R = 8% --- turn to a decimal = 0.08 so we have I and P and R....we need to find T now we sub I = PRT 960 = (6000)(0.08)(T) 960 = 480T 960/480 = T 2 = T check.. 960 = (6000)(0.08)(2) 960 = 960 (correct) so it took 2 years.....T = 2